OpenStudy (anonymous):

Calculate d/dx the equation is y=3x-9/2x+8

OpenStudy (anonymous):

Ok alecv write the derivitives of f(x) and g(x)

OpenStudy (anonymous):

ok f'(x) is 3 and g'(x) is 2

OpenStudy (anonymous):

Ok now remember the quotient rule: \[\frac{ f'(x)g(x)-f(x)g'(x) }{ g(x)^2 }\] Plug in your equations

OpenStudy (anonymous):

ok so I got 6/(2x+8)(2x+8)

OpenStudy (anonymous):

ok im checking your work

OpenStudy (anonymous):

I believe you forgot to distribute the minus sign again on the top: \[\frac{ 6x+24 - (6x-18) }{ (2x+8)^2 }\]

OpenStudy (anonymous):

hmm I tried that as well I believe it gives 42 it doesnt seem to like that either

myininaya (myininaya):

Matt is right and 42 is the numerator.

OpenStudy (anonymous):

so it should be 42/(2x+8)(2x+8)

myininaya (myininaya):

Maybe it wants you to simplify though.

myininaya (myininaya):

Do you know how to simplify?

OpenStudy (anonymous):

could you refresh my memory? haha x)

myininaya (myininaya):

2x+8 can be factored correct?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Well 42 and (2x+8) share a common factor

myininaya (myininaya):

The fraction can therefore be reduced.

OpenStudy (anonymous):

the common factor is 2 so it turns the numerator into 21

OpenStudy (anonymous):

does that then mean that the bottom becomes (x+4)?

myininaya (myininaya):

Well you have 2x+8=2(x+4) And you have (2x+8)^2 on bottom so that means you have 2(x+4)*2(x+4) Only one of those 2's get canceled with a 2 from the top so yeah you can say the numerator is 21 now but that would make the denominator what?

OpenStudy (anonymous):

ok so you got 42 over 2(x+4)*2(x+4) like you said one of the 2's has cancels out with the 42 making 21 so that leaves you with (x+4) times 2(x+4) right?

OpenStudy (anonymous):

cancelled*

myininaya (myininaya):

Yep Or 21/[2(x+4)^2]

OpenStudy (anonymous):

Thank you soo much! I appreciate the help! i'm sure i'll be asking for your help again at some point today my professor gave us a boat load of homework

OpenStudy (anonymous):

ok so again this time its asking for d/dx y=(x+6)(x^2-6) the x^2 is throwing me off